Solve by factoring: 4m^2 + 7m = 0
4m+7=0
4m=-7
m = -7/4
m = (-7/4, 0)
Where am I going wrong? What should I be doing? It was marked wrong with no explanation.

Answer by mananth(16949)

4m^2 + 7m = 0
m(4m+7)=0
if m= 0 the equation is satisfied.
if m=-7/4 the equation is satisfied.
so m= 0 OR -7/4
You are right but you must keep the m after factoring outside